hans is 3 years younger than his brother emil. in 8 years the sum of their ages will be 61. how old is hans
Answer by mathsmiles(68)

We need 2 variables, one for Hans (H) age and one for his brother Emil (E) age.

H = E-3. Hans is 3 yrs younger than Emil.

In 8 years, Hans will be H+8 and Emil will be E+8.
At that time, the sum of their ages will be 61.
(H+8) + (E+8) = 61

Now substitute (E-3) from above for H in the above equation:
((E-3)+8) + (E+8) = 61
Since this is all addition, we can remove the parens without any harm.
E - 3 + 8 + E + 8 = 61

Using the commutative law, let's re-order the terms:
E + E - 3 + 8 + 8 = 61
2E + 13 = 61
Subtracting 13 from both sides:
2E = 48
Dividing by 2 on both sides:
E = 24
Emil is 24; his brother is 3 years younger, so is 21.

Let's check out answer:
In 8 years, Emil will be 24 + 8 = 32; Hans will be 21 + 8 = 29
Adding their ages gives
32 + 29 = 61
61 = 61 Correct!